Using the data from FRED, enter the values for the second quarters of 2018 and 1987 for the M1 and M2 velocity of money: (Enter your responses exactly as they appear in FRED) M1 Velocity M2 Velocity Year-Month-Day (M1V) 2018-04-01 5.581 1987-04-01 6.457 1.734 M2V) 1.454 Using the data above and the values for nominal GDP below, calculate the value of the money supply for M1 and M2: (Enter your responses rounded to one decimal place.) M2 Money Nominal GDP (billions) M1 Money Supply (billions) Supply (billions) Year-Month-Day 2018-04-01 $20411.924S 1987-04-01 4806.16
Answer #1

Ans. We know that , Money Velocity × Money Supply = Nominal GDP.

So in 2018 , Money Supply ( M1) = $20411.924 ÷ 5.581 = $3657.39

So in 1987 , Money Supply ( M2 ) = $20411.924 ÷ 1.454 = $14038.46

In 2018 , M2 = $4806.16 ÷ 6.457 = $744.33

In 1987 , M2 = $4806.16 ÷ 1.734 = $2771.71

Best of Luck !! Thank You !!

A Thumsup would be Highly Appreciated !!

!!